OVHCloud Network Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#8310 — BHS <> MTL
Incident Report for Network & Infrastructure
Resolved
The 10*10G between Beauharnois and Montréal are down.

Right now it's going through the 10*10G between Newark and Beauharnois and a backup link.

Update(s):

Date: 2013-03-27 04:14:46 UTC
Montreal<>BHS is now up and running.

Everything is back online.


Date: 2013-03-27 04:10:05 UTC
Newark<>BHS link are now up and running.

Still waiting on Monreal<>BHS link. Should be quick.


Date: 2013-03-27 02:36:34 UTC
The field team is informing us that their new ETA is midnight, Montreal localtime.

Date: 2013-03-27 01:36:29 UTC
The last ETA we received: 10:30pm

Date: 2013-03-26 23:03:03 UTC
The repair (quite challenging because under a bridge) is still in progress. New ETA: 9pm, Montreal localtime.

Date: 2013-03-26 20:47:25 UTC
Repair operations are in progress. The impacted links should come back up after 6pm Montreal localtime.

Date: 2013-03-26 16:11:34 UTC
The fiber cut location has been identified. More than 400 fibers are impacted so the corrective splicing will take some time.

Date: 2013-03-26 15:55:49 UTC
Specialized teams have been dispatched along the fiber path to identify the cut location and proceed with repairs.

Date: 2013-03-26 13:55:28 UTC
We are investigating.
Posted Mar 26, 2013 - 13:55 UTC